OpenStudy (anonymous):

I understand the first part of the fundemantal threom of calculus, but the second part is killing, anybody got any tips or tricks to understand it?

OpenStudy (anonymous):

study

OpenStudy (anonymous):

Be more specific.

OpenStudy (anonymous):

can somone help me do a problem

OpenStudy (anonymous):

kinda coach me through it

OpenStudy (anonymous):

how would we begin this problem?

jimthompson5910 (jim_thompson5910):

The second part of the FTC (fundamental theorem of calculus) says that \[\int_{a}^{b} f(x)dx = g(b)-g(a)\] where g(x) is the antiderivative of f(x) In this case, \[\int (t+2)dt=\frac{1}{2}t^2+2t+C\] So \[\int_{0}^{x} (t+2)dt=(\frac{1}{2}x^2+2x+C)-(\frac{1}{2}(0)^2+2(0)+C)=\frac{1}{2}x^2+2x\]

jimthompson5910 (jim_thompson5910):

I'm just realizing that some books present the FTC in different order. Let me know if this is the portion you're referring to.

jimthompson5910 (jim_thompson5910):

Assuming you mean F(x) in the attachment, you have the right answer.